please help me find solution to this problem. Good day and thank you!
Troy made a rectangular pen for his dog using the side of the garage for one side and 30 feet of fencing for the other 3 sides. if the area is enclosed by the fence was 112 square feet, find the dimensions of the pen.

Answer by Alan3354(69443)

Troy made a rectangular pen for his dog using the side of the garage for one side and 30 feet of fencing for the other 3 sides. if the area is enclosed by the fence was 112 square feet, find the dimensions of the pen.
---------------------------
Area = L*W = 112
2W + L = 30 --> L = 30 - 2W
----
W*(30 - 2W) = 112
2W^2 - 30W + 112 = 0
W^2 - 15W + 56 = 0
W = 7, L = 16
-------------------
W = 8, L = 14